### v1

@@ 6 Conclusions
Due to the recent development in Lee metric code-based cryptography, in particular the lattice-based attack on the NIST submission FuLeeca, we analyzed the connection of Lee metric code-based cryptosystems and the corresponding lattice problems in the $\ell_{1}$- and the $\ell_{2}$-norm.
@@
These results show that when (Hamming and) lattice techniques can be used to break Lee metric code-based cryptosystems. Hence, this can tell us which parameters should be avoided when designing new public key encryption schemes or digital signatures using Lee metric error correcting codes. In particular, when $q$ is chosen extremely small, then Hamming-based coding techniques can be used to attack Lee metric cryptographic schemes. On the other hand, for large $q$, $\ell_{1}$-lattice techniques might be applicable, however, using $\ell_{2}$-techniques will most likely not work. Moreover, when $q$ is very large, then the idea of the FuLeakage attack of using the lattice generated by a basis of the code (instead of the whole $\mathsf{Construction\leavevmode\nobreak\ A}$ lattice) has a higher success probability than for smaller $q$. This means that in general, $q$ should be chosen large enough such that Hamming attacks are not applicable but also small enough that a large part of the $\mathcal{L}_{\mathsf{A}_{\mathbf{G}}}$ lattice is not contained in $\mathcal{L}_{\mathsf{A}}$. Furthermore, $n$ should be chosen large enough that the $\ell_{2}$-norm is not a good approximation of the $\ell_{1}$-norm. The exact parameters depend on the actual cryptosystem, the relationship of modulus, length, and minimum distance of the code and need to be investigated when designing a new Lee metric code-based cryptosystem.
@@
In order to prove Theorem 2.19, we use Blichfeldt’s theorem on non-full dimensional lattices. In the following, our proofs are based on the proofs from [10, Theorem 20-21] and [14, Theorem 5-6].

### v2

@@ 6 Conclusions
Due to the recent developments in Lee metric code-based cryptography, in particular the lattice-based attack on the NIST submission FuLeeca, we analyzed the connection of Lee metric code-based cryptosystems and the corresponding lattice problems in the $\ell_{1}$- and the $\ell_{2}$-norm.
@@
These results show when (Hamming and) lattice techniques can be used to break Lee metric code-based cryptosystems. Hence, this can tell us which parameters should be avoided when designing new public key encryption schemes or digital signatures using Lee metric error correcting codes. In particular, when $q$ is chosen extremely small, then Hamming-based coding techniques can be used to attack Lee metric cryptographic schemes. On the other hand, for large $q$, $\ell_{1}$-lattice techniques might be applicable, however, using $\ell_{2}$-techniques will most likely not work. Moreover, when $q$ is very large, then the idea of the FuLeakage attack of using the lattice generated by a basis of the code (instead of the whole $\mathsf{Construction\leavevmode\nobreak\ A}$ lattice) has a higher success probability than for smaller $q$. This means that in general, $q$ should be chosen large enough such that Hamming attacks are not applicable but also small enough that a large part of the $\mathcal{L}_{\mathsf{A}_{\mathbf{G}}}$ lattice is not contained in $\mathcal{L}_{\mathsf{A}}$. Furthermore, $n$ should be chosen large enough that the $\ell_{2}$-norm is not a good approximation of the $\ell_{1}$-norm. The exact parameters depend on the actual cryptosystem, the relationship of modulus, length, and minimum distance of the code and need to be investigated when designing a new Lee metric code-based cryptosystem.
Acknowledgements
Carlos Vela Cabello is supported by the Grundlagenforschungsfond (GFF) of the University of St.Gallen, project no. 2260780. The authors wish to thank Cameron Foreman and Kevin Milner for useful comments on the manuscriptreviewing the manuscript.
@@ Acknowledgements
In order to prove Theorem 2.19, we use Blichfeldt’s theorem on non-full dimensional lattices. In the following, our proofs are based on the proofs from [12, Theorem 20-21] and [17, Theorem 5-6].